Project

General

Profile

News

Update 24 September 2016

Added by Kyberd, Paul 12 months ago

Please join me in congratulating all involved in bringing the MICE magnets to current in a configuration typical of the settings required to deliver the field-on scattering programme at Step IV. The magnets have been at current since the evening of Thursday 22Sep16 and a full 48-hour soak test is now in progress. On Friday 23Sep16, the the first data was taken in this configuration with 140 MeV and 170 beams, i.e. we’ve started the empty-channel data taking for the field-on scattering programme.

Update 22 August 2016

Added by Kyberd, Paul 12 months ago

Success at the summer conferences:
MICE presentations at the summer conferences, ICHEP and NuFact, were based around the papers that are in preparation. The bases of the presentations were:
• The first reconstruction of emittance using a beam sampled one particle at a time. These data validate the reconstruction technique we plan to use to study cooling;
• The first measurement of multiple Coulomb scattering (MCS) of muons on lithium-hydride (LiH) using field-off data collected earlier this year; and
• The final results of the optimisation of the cooling-demonstration lattice.

The quality of the presentations was uniformly high. Steve Boyd made the parallel-session presentation at ICHEP, while posters were presented by Victoria Blackmore and Tanaz Mohayai. At NuFact the three parallel-session presenters were Francois Drielsma, Ryan Bayes and Yordan Karadzhov. Each of the presenters gave a very professional impression, reflecting well on the months of hard-work that led to the excellent new results.

Successful combined test of SSU and focus coil:
On the 02Sep16 the magnet-string commissioning passed a most significant milestone. The combination SSU was run in combination with the FC at currents representative of those that are required to deliver the solenoid-mode scattering programme. The field in the tracking volume was 3T, the operating point chosen for the Step IV programme.

SSU was run up so that the ECE combination carried 206A, M2 190A and M1 100A. Then, the FC was run to 75A in small current steps. The attached plot shows the ramp of the FC with SSU energised. The blue trace is the FC current, the black trace is the force measured by the load cells on the FC and the magenta trace is the inter-module force calculated using the currents in the various windings and the coil dimensions. This success will allow the magnet and operations teams to plan the preparation for the User Cycle with confidence.

Magnet-string commissioning will continue next week and on to the start of the next ISIS User Cycle on 13Sep16.

Data in the bank:
A substantial quantity of high-quality data with the superconducting-magnet channel in various configurations was taken in the most recent ISIS User Cycle. This data can be used to study the alignment of the magnetic axis of each magnet individually and to make initial studies of the solenoid-mode optics. It is important that these studies are documented and published the results for future reference.

Commissioning and Hall status update:
Progress has been made in identifying issues in the modules that control the “trim current” to the end coils in the solenoids. A test remains to be performed to demonstrate the stability of the end coils in operation. The source of transients observed in the voltage-tap signals of SSU and SSD remains to be understood fully. In the last User Cycle, stable operation of SSU and SSD was achieved with a voltage-tap threshold of 200mV. A control-software update for the spectrometer solenoids has been prepared at Fermilab and will be deployed next week. Further tests to elucidate the source of the transient signals will then be performed.

There has been steady progress on the preparation of the liquid-hydrogen vessel in R9. The absorber bodies have been passed "good to go" and the MLI has been fitted. Pipe-work modifications are complete. Final level-sensor checks have been carried out and tape to reduce the emissivity of the warm surfaces of the magnet have been installed. The end-caps have been mounted on cradles to allow final preparations to be made. The end caps will be fitted towards the end of next week. Evacuation of the module and the "interspace" can then begin.

The upgrade to the Hall water system is now complete; the demineralised and “inhibited” circuits are now separate; the final modifications having been made on Thursday 01sep16.

The Kaiser compressor serving the Linde refrigerator that cools the decay solenoid was serviced. Following the service, air contamination of the helium was observed and a number of small leaks were found. Kaiser engineers made a visit to the Hall on 31Aug16 and identified two leaks. Work continues to establish that the system is leak tight and, if it is not, to address the remaining issues.

The remedial work on the air-conditioning units is complete. The Hall now has five functioning air conditioners. This, together with the decreasing air temperatures as autumn comes to Oxfordshire, is estimated to be sufficient to stabilise the ambient temperature in the Hall.

ISIS User Cycle 2016/03 (13Sep16 — 29Oct16):
Planning for the next User Cycle is now well underway. The run plan includes data-taking to allow detailed studies of the optics of the Step IV solenoid-mode lattice. The priority for the Cycle is empty-absorber data-taking for the Step IV, field-on scattering programme.

Shift sign-up in CHEESE is live and all are encouraged to sign up.

Update 22 July 2016

Added by Kyberd, Paul 12 months ago

Dear All,

Over the past week a number of milestones were passed in the commissioning of the superconducting magnets.

The focus coil has been in routine operation in solenoid mode for data-taking. Data has been taken to validate the magnetic alignment of the module. Analysis is underway.

The end and centre coils of both solenoids have been operated at a current of 140A. Transients were observed on the quench-detection signals of both the upstream and the downstream solenoids. The upstream solenoid (SSU) exhibits a higher rate of such transients than the downstream solenoid (SSD), the absolute size of the excursions is small. After a number of investigations were carried out, including the study of the movement of the partial return yoke, the relative movement of the yoke and the solenoids and the movement of the solenoids themselves. When all of these investigations failed to identify clearly the source of the transients, it was decided to increase the trip threshold to 200mV and run the magnets individually for a period to observe the rate of transients and whether there was any trend for the transients to grow. These tests were successful in that the transients never exceeded the 200mV threshold and were observed to be of relatively constant amplitude.

Having run both magnets individually to allow data taking for magnetic alignment, on Wednesday 20Jul16, both SSU and SSD were brought to 140A. In the combined combination SSU + SSD, the solenoids were observed to behave just as they did in stand-alone running. Stable operation was established and data was taken with more than one superconducting solenoid in operation for the first time.

The attached slides show a selection of plots from the data taking on Wednesday 20Jul16; I hope the slides are self explanatory. The distribution of space points in the upstream and downstream trackers appears to follow the beam envelope predicted by Chris Roger’s linear-optics calculation. The phase-space distributions obtained using the reconstructed tracks are also shown. The result of a MAUS simulation of the conditions under which data was taken on the 20Jun16 is shown in the slides for comparison. The simulation gives a fair, qualitative, description of the data. There is, of course, more work to do to understand the differences in detail.

The results presented in the slides are the result of the hard work of all; a really important step has been made!

Moving forward, the magnet and operations teams are discussing how best to approach the goal of taking data with all superconducting magnets—in the configuration SSU+FC+SSD before the end of the present User Cycle.

Looking forward to discussing the progress and results at CM45!

Ken

Update 01 July 2016

Added by Kyberd, Paul 12 months ago

Dear All,

This is just a brief status update now that ISIS Cycle 2016/02 is under way and magnet commissioning is in full swing.

Please do try and make it to the “special VC”, today at 15:30, at which Durga and Steve Boyd will rehearse their talks for Neutrino 2016.

With best wishes …

Ken

Status update:

Physics and analysis:
• The next of our regular video conferences will be on Thursday 14Jul16.  The meeting will be devoted to presentations from the authors of the three papers that are presently in preparation.  In each case, the results that the authors seek to make public, i.e. the results that they request be given “Preliminary” status will be presented.  These result will then be presented at ICHEP and at the summer conferences that follow.
Step IV:
• Focus-coil module:
◦ Over the past week or so, the focus coil has been brought into operation.  The magnet was operated successfully in solenoid mode and ran to 90A without quenching.  Last night a 12-hour soak test at 90A was successfully performed, completing the standalone commissioning of the module.
• Spectrometer solenoids:
◦ The check-out of the upgraded quench-detection and quench-protection systems has been completed successfully.  Current has been run in both spectrometers.  The final steps in the standalone commissioning are now being carried out.  It is anticipated that the standalone commissioning will be completed by Sunday 03Jul16.  The next step will be to run both SSU and SSD at the same time.  If all goes well, combined operation of SSU and SSD is scheduled for Sunday 03Jul16.
• Commissioning and operation for data taking: 
◦ Once combined operation of SSU and SSD has been established, commissioning of the MICE superconducting-magnet channel will begin.  This will start on Monday 04Jul16.  A general scheme for this activity has been agreed.  The details of the programme will be discussed this afternoon (Friday 01Jul16).  Steve Boyd will disseminate the operations plans once they have been finalised.

Update 03 June 2016

Added by Kyberd, Paul 12 months ago

Dear All,

At the last Video Conference we discussed a number of descoped configurations by which we could deliver a demonstration of ionization cooling while substantially reducing the cost and risk of the programme to the STFC. These options had been prepared in anticipation that MICE-UK would be asked to submit a proposal for the resources to implement a descoped cooling demonstration.

On the same day, the STFC Executive Board met; the future of the MICE programme was one of the items that the Board considered. On Wednesday 01Jun16, in an email to the MICE Project Board in which I was included, I received notification of the outcome of the STFC Executive Board’s decision. The Board has decided “… that MICE should only be funded for extended Step IV operations that will allow a full set of measurements of the phase space, which should be completed by mid-2017.”

The consequences of this decision are that we shall be able to carry out the full Step IV programme defined in the run plan we prepared for the November 2014 MPB meeting. However, the MICE-UK collaboration will not now be invited to submit a proposal to carry out a descoped cooling demonstration.

The excellence of the work of the collaboration was noted in the same communication which contains the comments:
• “Research that is truly at the frontier is never guaranteed to run smoothly and it is a huge credit to scientists and engineers who have been working on the experiment at RAL that they have managed to push the envelope to beyond the limit of current technology. We also recognise the huge effort that the international team has put into MICE over the past decade and encourage all to extract as much science as possible over the coming months from this extended running in the Step IV configuration.“

Earlier this week, Colin and I visited the cavity preparation area at LBNL.  Progress is impressive and the single-cavity modules are on track to be delivered to RAL over the summer.  At the same time, the work on the RF amplifiers is going forward in the UK.  We therefore have all the elements to discuss how we might extend the programme to include the study of RF in the magnetic field of our channel.

Clearly this is not the outcome for which we had hoped. However, we do have the opportunity to carry out the full Step IV programme. The collaboration has demonstrated its ingenuity and exceptional ability to find innovative solutions to delivering a demonstration of the ionization-cooling technique. We should now return to the Step IV run plan and seek a programme by which each of the factors critical to the performance of an ionization-cooling channel are determined.

Our regular video conference is scheduled for 09Jun16 and I will seek further clarifications on the future development of our programme in time for this meeting. In the mean time, we must remain focused on completing the three papers, two on Step IV data, that are in preparation and continue to commission the magnets on the beam line in preparation for the June/July ISIS User Cycle.

If you would like to discuss the situation with me, please don’t hesitate to call my work mobile (the number is in the signature field below).

With best wishes …

Update 21 May 2016

Added by Kyberd, Paul 12 months ago

Dear All,

A lot of water has passed under the bridge in the couple of weeks since the MICE-UK cost to completion (CtC) review took place on the 26Apr16.

As I reported at the 05May16 VC (https://indico.cern.ch/event/525339/), the CtC panel recommended that the scientific output of Step IV be maximised. Therefore, should a cooling demonstration programme not go forward, the CtC panel recommended that Step IV running be extended such that the full programme defined at the Nov14 meeting of the MICE Project Board (http://micewww.pp.rl.ac.uk/projects/governance/wiki/2014-11-24-Review) be completed. Following this recommendation, STFC is now preparing to put in place the resources necessary to allow MICE to run until the end of the summer 2017 ISIS User cycle (Aug17) and to support the subsequent analysis activity.

The CtC panel also recommended that we find a configuration by which the cooling demonstration can be executed that has a cost and risk profile substantially lower than that reported to the cost-to-completion review. The framework for the development of the new configuration as well as the constraints on the eventual programme were outlined at the 05May16 VC and summarised in http://micewww.pp.rl.ac.uk/attachments/6188/2016-05-04-Descope.pdf. Work along these lines has been proceeding rapidly over the past couple of weeks.

At very short notice, I was asked by STFC for a briefing on the progress being made on the identification of low-risk, low-cost (descope) option for the cooling demonstration. Colin and I prepared a short memorandum in response; it is posted at: http://micewww.pp.rl.ac.uk/documents/177. The memo is to be used as input to a paper that is being prepared for submission to the STFC Executive Board (EB). The STFC EB meets next Thursday (26May16) and will consider whether MICE-UK will be invited to bring forward a proposal for the execution of the cooling demonstration.

Much work has been done under considerable pressure by many members of the collaboration. It is now time to take stock of the excellent work that has been done and to discuss the work that is still to come. I’ve therefore arranged an “extraordinary” video conference for:

•     Thursday 26May16 at the usual time (15:00 BST)
A WWW page has been prepared at:
https://indico.cern.ch/event/533773/

I understand that this time may not convenient for all; in fact the UK university community is taking industrial action next Wednesday and Thursday. However, I believe it to be important to share the information and to broaden the discussion and propose therefore to hold the meeting as defined above.

Appended is one of my periodic updates on progress.

With best wishes …

Ken

Status update:

Physics and analysis:
An analysis workshop was held on 19May16 (see http://micewww.pp.rl.ac.uk/projects/analysis/wiki/FF-2016-05-19-workshop).

A first meeting between the wise-people and the authors of the emittance paper took place on the 18May16. The meeting was a success, with useful comments on the analysis and the presentation of the information in the MICE Note. The analysis workshop received contributions on the various outstanding systematic studies that are required to tie down the systematic error on the emittance measurement. The present status of the emittance measurement is summarised in the first attachment (taken from V. Blackmore’s talk at the 05May16 VC). Work continues to understand and resolve remaining differences between the Monte Carlo and the data and the remaining systematic issues.

Strides have also been made in the preparation of the scattering measurement. I’ve picked the slide that R. Bayes presented on the results of his analysis in the 3-200 setting as an example (second attachment). Work continues to understand systematic effects and to resolve details in the analysis. The analysis is now sufficiently advanced for wise-people to be appointed so that the results can be prepared efficiently for publication.

The workshop also received presentations on the status of the work on the reconfiguration of the cooling demonstration and discussed at length the pros and cons of the various options. Please do look at the presentations to see how much has been achieved thus far.

It is now important that the cooling-demo design paper is published as soon as possible. It had been the intention to include the reconstruction performance in the paper to document the precision of the experiment. While progress has been made in understanding differences in the upstream and downstream emittance measurement, we now need to decide whether to limit the scope of the paper to the optics and cooling performance of the lattice, leaving the details of the simulated performance of the reconstruction to a subsequent publication.

Step IV:
Steady progress is being made on finalising the details of the Step IV:
• The h/w for the revised QD/QP system for the spectrometer solenoids is now in place. The power supplies have been operated into shorts and the system is now ready for tests to continue next. The commissioning of the system with current in the magnets will then follow;
• Preparations to separate the demineralised water circuit from the “inhibited” water circuit continue. Contractual issues have caused the work to go a little more slowly than we had planned. Therefore contingency plans are being put in place to ensure that the conventional magnets and the superconducting magnets can be run simultaneously in the June/July ISIS User cycle; and
• Work continues to schedule on the LH2 system. The absorber vessel has been “passed fit”, the procurement of the FC-module end flanges is underway and necessary wiring modifications are being carried out in R9. Work on the chiller required for the water-system reconfiguration has displaced work on the H2 gas supply; this work will now be done during the June/July ISIS User Cycle.

Cooling demonstration:
• The components for the PRY are complete and are on their way to RAL.
• The preparation of the single-cavity modules and the RF-power amplifiers proceeds to schedule.

Update 11 April 2016

Added by Kyberd, Paul 12 months ago

Dear All,

The reviews of last week, together with the presentations the week before at CM44, provide an excellent summary of the status of the project.

I’m sending this brief update only to note the increments of progress made over the past week and to invite you all to the next “MICE briefing” which will be held in R9 on Thursday the 14th April 2016. The purpose of this briefing is to celebrate the hard work that allowed such an excellent impression to be given in the recent reviews and to allow us to discuss informally the issues raised by the revised framework for the recovery of the downstream solenoid.

With best wishes …

Ken

MICE briefing:
• R9; 15:00 on Thursday the 14th April 2016

Status update:

Physics and analysis:
The status of the analysis is succinctly summarised in the presentations to the MPB. The principal advancement over the past week is from Francois Drielsma who has identified issues in the handling of the magnetic field maps in the Monte Carlo. The issues relate to the geometrical “placement” of the maps. When these issues are corrected, the data/Monte Carlo comparison plots for the trackers and the TOFs are much improved. Further tuning will continue, but I am optimistic that the work for the emittance-reconstruction paper can now proceed with confidence.

We shall need to be careful to plan work on the emittance measurement and the scattering measurement such that we can prepare the preliminary results in time for the summer conferences. A practical hard deadline is afforded by the ICHEP 2016 paper-submission deadline which has just been extended to 31May16. Chris Rogers will be organising the analysis meetings such that the steps necessary for results to gain “Preliminary” status are carried out in a timely manner.

Step IV:
Of course, the hardware status is carefully summarised in the RLSR/MPB presentations. Progress beyond that noted there includes:
• The LiH absorber has been extracted from the focus-coil and moved to safe storage. Electrical work on the FC itself continues in preparation for the cool-down of the module later this month, which will be followed by stand-alone commissioning;
• Preparations to separate the “demin” and “inhibited” water circuits is underway. A chiller, loaned to us by ISIS, will be used to ensure a minimum of disruption to the circuit that cools the compressors serving the upstream and downstream solenoids.
• The remedial work on the water system has been given priority as the risk to the compressors of continued us of demineralised water is too high. This has the necessary consequence that weekend running in April 2016 has been cancelled. The weekend data-taking would have been used to “top up” the empty-channel data sets for the LiH scattering measurement. These data sets appear to be reasonably well populated and it seems likely that they will be sufficient to allow the scattering distributions to be unfolded.
• The installation of components of the QD and QP systems for the spectrometer solenoids continues. The necessary modifications to Rack Room 2 are underway, the DCCTs have been installed and the framework that will support the external dump resistors is being assembled in R9.

Future meetings:
• Video conferences:
• 05May16
• 09Jun16
• 14Jul16
• Physics workshop:
• 28th April 2016 at RAL

Update 01 February 2016

Added by Kyberd, Paul 12 months ago

Status update:

Step IV:
Work in the MICE Hall is focussed on the necessary preparations for the LiH scattering measurements that are planned for the next ISIS User Cycle.  It has been decided to adjust the positions of the spectrometer solenoids to account for the measured offset between the magnetic and mechanical axes.  Before this work can be undertaken, appropriate bellows to join the spectrometer solenoids to the focus-coil module must be procured.  

Many “house-keeping” tasks have been completed (e.g. flushing of the beam-line magnets, DKsoln fridge and compressor servicing, implementation of “pseudo-kinematic” locators for TOF2, KL and EMR …).  

Work in R9 is focussed on preparing for the extraction of the liquid-hydrogen vessel so that the leaking seal can be repaired.   

A new quench-detection/quench-protection system for the spectrometer solenoids has been devised by the spectrometer-solenoid team.  Detailing of the specification of the various components is under way and measurements of critical parameters will be made in Rack Room 2 this week.  The spectrometer-solenoid recovery plan continues to be developed.

Software and analysis:
The systematic check-out of the tracker software (Dobbs, Heidt, Hunt, Overton, Uchida) is reaching the end of the first pass through the code.  The sign-flip issues in the reconstruction have been addressed and the algorithm check-out from cluster finding, through pattern recognition to the track fit has been completed successfully.  Features in the tracker simulation have also been understood and addressed.  A. Dobbs is preparing a new release of MAUS that will contain fixes to all known issues.  With this release, D. Rajaram will make a consistent reprocessing of all the data that was taken in 2015.

When the data is reprocessed, the alignment constants derived by F. Drielsma will be included.  The first attachment shows the status of the alignment of the instrumentation from TOF1, through both the trackers to TOF2, KL and EMR.  

The analysis of the data taken on the 7th October 2015 with the end and centre coils of the upstream solenoid at nominal current is progressing.  The second attachment (Blackmore) shows the emittance reconstructed from tracks in the upstream tracker plotted in 5 MeV/c momentum bins from 160 MeV/c to 245 MeV/c.  The data has been split into two samples of equal size.  The first is used to calculate the emittance at the entrance to the upstream tracker (station 5, black solid circles).  The second is used to calculate the emittance at the exit of the tracker (station 1, red solid triangles).  While more work needs to be done, the plot indicates that we shall be able to make a good measurement of emittance using the single-particle technique.

The xenon scattering data taken just before Christmas is also being analysed.  The third attachment shows the scattering angle determined by taking the dot product between the track direction determined in the upstream tracker with that determined downstream (Nugent).  The scattering analysis needs to include the alignment corrections outlined above, but, the preliminary analysis indicates that it will be possible to unfold the scattering distribution.

Video conference:
The next video conference will take place on Thursday 04Feb16.  

Update 15 January 2016

Added by Kyberd, Paul 12 months ago

Dear All,

The “big” event this week was the first formal review of the Optics Analysis by an external panel of experts. This was both a success and an important step for the collaboration.

The excellence of the material, which is collected on the review page (http://micewww.pp.rl.ac.uk/projects/governance/wiki/2016-01-14-15-OpticsReview), speaks for itself. The document prepared to brief the reviewers (Status of the MICE Optics and Analysis rev 1), the presentations and the material prepared in response to the reviewers’ questions will be an important resource in the weeks to come.

I would like to congratulate Chris Rogers and the analysers not only for the excellent talks but also for demonstrating an impressive grasp of the accelerator, detector and experimental physics of MICE and of the outstanding issues that need to be addressed; their command of detail, apparently effortless, was was recognised by the reviewers.

The two days of the review were both enjoyable and inspiring; recording our achievements to date, the capabilities we have at Step IV and the potential for impact with the demonstration of ionization cooling.

A brief status report follows …

With very best wishes for 2016 …

Ken

Status update:

Magnets:
• Absorber/focus-coil module:
⁃ A leak, perhaps the leak, in the hydrogen system has been identified. It has been located in the indium seal of the downstream window. This is excellent news. The module now needs to be taken to R9 for the hydrogen vessel to be withdrawn and a repair effected. The move of the AFC from the Hall to R9 is scheduled to take place on Friday 22Jan16.
• Spectrometer solenoids:
⁃ Progress has been made on the design and specification of the QD/QP system required to allow operation of SSU and SSD at Step IV. Over the coming two weeks the system will be specified. Discussions about the implementation will then gather momentum.
⁃ The SSD repair plan is being developed. Clearly a difficult problem. A joined up approach is being taken with potential manufacturers being considered on both sides of the Atlantic in a coherent effort between US personnel and personnel at RAL.
• Decay solenoid:
⁃ The repaired FUG power supply is still expected to be returned to RAL by the end of February 2016. Once receipted, the module will need to be installed and commissioned.
⁃ The Linde fridge has been serviced. Next week, the helium compressor that serves the Linde fridge will be serviced.

MLCR:
• The controls and monitoring network re-configuration is in progress. This is necessary to separate the controls from the daq network and to increase the robustness of communications between the various critical modules on the controls network.

Update 1st February

Added by Kyberd, Paul over 2 years ago

Step IV:

Work in the MICE Hall is focussed on the necessary preparations for the LiH scattering measurements that are planned for the next ISIS User Cycle. It has been decided to adjust the positions of the spectrometer solenoids to account for the measured offset between the magnetic and mechanical axes. Before this work can be undertaken, appropriate bellows to join the spectrometer solenoids to the focus-coil module must be procured.

Many “house-keeping” tasks have been completed (e.g. flushing of the beam-line magnets, DKsoln fridge and compressor servicing, implementation of “pseudo-kinematic” locators for TOF2, KL and EMR …).

Work in R9 is focussed on preparing for the extraction of the liquid-hydrogen vessel so that the leaking seal can be repaired.

A new quench-detection/quench-protection system for the spectrometer solenoids has been devised by the spectrometer-solenoid team. Detailing of the specification of the various components is under way and measurements of critical parameters will be made in Rack Room 2 this week. The spectrometer-solenoid recovery plan continues to be developed.

Software and analysis:
The systematic check-out of the tracker software (Dobbs, Heidt, Hunt, Overton, Uchida) is reaching the end of the first pass through the code. The sign-flip issues in the reconstruction have been addressed and the algorithm check-out from cluster finding, through pattern recognition to the track fit has been completed successfully. Features in the tracker simulation have also been understood and addressed. A. Dobbs is preparing a new release of MAUS that will contain fixes to all known issues. With this release, D. Rajaram will make a consistent reprocessing of all the data that was taken in 2015.

When the data is reprocessed, the alignment constants derived by F. Drielsma will be included. The first link shows the status of the alignment of the instrumentation from TOF1, through both the trackers to TOF2, KL and EMR.

The analysis of the data taken on the 7th October 2015 with the end and centre coils of the upstream solenoid at nominal current is progressing. The second link (Blackmore) shows the emittance reconstructed from tracks in the upstream tracker plotted in 5 MeV/c momentum bins from 160 MeV/c to 245 MeV/c. The data has been split into two samples of equal size. The first is used to calculate the emittance at the entrance to the upstream tracker (station 5, black solid circles). The second is used to calculate the emittance at the exit of the tracker (station 1, red solid triangles). While more work needs to be done, the plot indicates that we shall be able to make a good measurement of emittance using the single-particle technique.

The xenon scattering data taken just before Christmas is also being analysed. The third attachment shows the scattering angle determined by taking the dot product between the track direction determined in the upstream tracker with that determined downstream (Nugent). The scattering analysis needs to include the alignment corrections outlined above, but, the preliminary analysis indicates that it will be possible to unfold the scattering distribution.

Video conference:
The next video conference will take place on Thursday 04Feb16.

http://micewww.pp.rl.ac.uk/images/2016-02-02_alignment.pdf
http://micewww.pp.rl.ac.uk/images/scattering_change-1.pdf
http://micewww.pp.rl.ac.uk/images/2016-02-02/run7469_5MeVPzSlices_Emittance.pdf

(11-20/114)

Also available in: Atom